### v2

# Deflection angle in the strong deflection limit for static axisymmetric spacetimes: local curvature, matter fields, and quasinormal modes

Abstract
We investigate the deflection of photons in the strong deflection limit within static, axisymmetric spacetimes possessing reflection symmetry. As the impact parameter approaches its critical value, the deflection angle exhibits a logarithmic divergence. This divergence is characterized by a logarithmic rate and a constant offset, which we express in terms of coordinate-invariant curvature evaluated at the unstable photon circular orbit. The curvature contribution is encoded in the electric part of the Weyl tensor, reflecting tidal effects, and the matter contribution is encoded in the Einstein tensor, capturing the influence of local energy and pressure. We also express these coefficients using Newman–Penrose scalars. By exploiting the relationship between the strong deflection limit and quasinormal modes, we derive a new expression for the quasinormal mode frequency in the eikonal limit in terms of the curvature scalars. Our results provide a unified and coordinate-invariant framework that connects observable lensing features and quasinormal modes to the local geometry and matter distribution near compact objects.
I Introduction
Recent observational advances have significantly enhanced our capacity to explore spacetime geometry in the strong-field regime. In particular, the imaging of black hole shadows has vividly illustrated how light propagates near compact objects [1, 2]. At the same time, gravitational lensing has long provided a theoretical basis for understanding the motion of photons through curved spacetimes (see, e.g., Ref. [3] for a review).
When light rays pass extremely close to a compact object, the deflection angle becomes large. In this regime, the strong deflection limit (SDL) provides a useful approximation for analyzing gravitational lensing. The deflection angle diverges logarithmically as the impact parameter approaches its critical value, with a characteristic rate and a constant offset, known as the SDL coefficients. A systematic formulation of this behavior was developed by Bozza [4], and has since been refined and extended in several works [5, 6]. These methods have been widely applied to static, spherically symmetric spacetimes [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], and have further inspired a variety of extensions and applications [25, 26, 27, 28]. Furthermore, this formulation is applicable to the reflection-symmetric plane of axisymmetric spacetimes [29, 30, 31, 32, 33, 34].
Recently, it has been shown that, in static, spherically symmetric spacetimes, the SDL coefficients can be expressed in terms of local geometric and matter field quantities at the photon sphere [35]. This coordinate-independent formulation refines earlier coordinate-based approaches and provides a framework that links local curvature and matter distribution to observable features of strong gravitational lensing. However, whether a similar local and coordinate-independent description remains valid in less symmetric spacetimes has not yet been fully clarified.
Independently of these developments, quasinormal mode (QNM) frequencies have also been found to reflect the properties of unstable photon circular orbits, particularly in the eikonal limit [36, 37]. Interestingly, a connection between the SDL and QNMs was pointed out in previous studies [38, 39], which showed that the imaginary part of the QNM frequency in the eikonal limit is inversely proportional to an SDL coefficient. More significantly, recent work [35] has shown that, in static and spherically symmetric spacetimes, the QNM frequency in the eikonal limit can be expressed in terms of local geometric and matter field quantities evaluated at the photon sphere. This coordinate-invariant formulation provides a physical insight into the QNM frequency as a manifestation of both spacetime curvature and matter distribution near the unstable photon orbit, thereby offering a unified geometric perspective complementary to that of the strong deflection limit.
The relation between QNMs and the SDL is particularly significant in light of recent gravitational wave observations, where the ringdown phase has been detected and matched with theoretical predictions [40, 41, 42]. This underscores the importance of understanding the local geometry near the photon sphere in interpreting strong-field gravitational wave signals.
In this work, we generalize the coordinate-invariant formulation of the deflection angle in the SDL to static and axisymmetric spacetimes with reflection symmetry. We show that the SDL coefficients can be expressed in terms of local geometric and matter field quantities evaluated at the unstable photon circular orbit. Furthermore, we reformulate these coefficients using curvature-based quantities, such as the electric part of the Weyl tensor and Newman–Penrose (NP) scalars and apply the formalism to several explicit spacetimes to illustrate its physical relevance. Using the relation between the SDL and the QNM, we derive a new expression for the QNM frequency in terms of local curvature and matter field quantities.
This paper is organized as follows. In Sec. II, we introduce the general formalism of photon dynamics in static, axisymmetric spacetimes and derive the conditions for the existence of circular photon orbits. In Sec. III, we derive the deflection angle in the SDL by isolating the logarithmically divergent contribution from the photon trajectory integral. In Sec. IV, we recast the SDL expression in a coordinate-invariant form, demonstrating that it depends only on the circumferential radius of the unstable photon circular orbit and local curvature. In Sec. V, we relate this expression to the matter field quantities. In Sec. VI, we further recast the deflection angle in terms of the NP scalars. In Sec. VII, we present several applications of the formalism developed herein. In Sec. VIII, we establish the correspondence between QNM frequencies and SDL coefficients in static, axisymmetric spacetimes. Finally, in Sec. IX, we summarize our findings and discuss their implications for gravitational lensing in the SDL.
Throughout this paper, we employ the abstract index notation [43] and use geometrized units in which the gravitational constant and the speed of light are set to unity.
II Unstable photon circular orbits in static, axisymmetric spacetimes
We consider a general static, axisymmetric spacetime, which admits Killing fields corresponding to time translations and rotations, adapted to the coordinates $t$ and $\varphi$, respectively. The canonical form of the line element is given by
$\mathrm{d}s^{2}=-e^{2\>\!\psi}\,\mathrm{d}t^{2}+e^{-2\>\!\psi}\left[\>\!e^{2\>\!\gamma}\left(\mathrm{d}\rho^{2}+\mathrm{d}\zeta^{2}\right)+W^{2}\>\!\mathrm{d}\varphi^{2}\>\!\right],$ (1)
where $\psi$, $\gamma$, and $W$ are functions of $\rho$ and $\zeta$ [44]. We assume a reflection symmetry (i.e., a $\mathbb{Z}_{2}$ symmetry) with respect to the $\zeta=0$ plane, which requires that $\psi$, $\gamma$, and $W$ are even functions of $\zeta$:
$\psi(\rho,-\zeta)=\psi(\rho,\zeta),\quad\gamma(\rho,-\zeta)=\gamma(\rho,\zeta),\quad W(\rho,-\zeta)=W(\rho,\zeta).$ (2)
We define the auxiliary function
$R(\rho,\zeta)=e^{-\psi}W,$ (3)
which represents the circumferential radius around the symmetry axis ($\rho=0$). We assume $\rho\in(0,\rho_{\infty})$ and that $R$ diverges as $\rho\to\rho_{\infty}$ for any fixed $\zeta$. Furthermore, on the $\zeta=0$ plane, we assume that the spacetime is asymptotically locally flat; that is, as $\rho\to\rho_{\infty}$, the metric functions approach their Minkowski values: $e^{2\psi}\to 1$, $e^{2\gamma}\to 1$, and $W\to\rho$.
We assume that photons follow null geodesics. Restricting our analysis to the reflection-symmetric plane $\zeta=0$, where every initially confined photon remains there, we obtain the reduced Lagrangian
$\mathscr{L}=\frac{1}{2}\left[\>\!-e^{2\>\!\psi}\>\!\dot{t}^{2}+e^{-2\>\!\psi}\left(e^{2\>\!\gamma}\>\!\dot{\rho}^{2}+W^{2}\dot{\varphi}^{2}\right)\>\!\right],$ (4)
where $\psi$, $\gamma$, and $W$ are evaluated at $\zeta=0$, and the overdot denotes differentiation with respect to an affine parameter along the null geodesic. Since $\mathscr{L}$ is independent of $t$ and $\varphi$, the corresponding conjugate energy and angular momentum,
$E$ $=e^{2\>\!\psi}\>\!\dot{t},$ (5)
$L$ $=e^{-2\>\!\psi}W^{2}\dot{\varphi},$ (6)
respectively, are conserved.
By combining the null condition $\mathscr{L}=0$ with Eqs. (5) and (6), we obtain
$\dot{\rho}^{2}+e^{-2\>\!\gamma}\left(\frac{e^{4\>\!\psi}}{W^{2}}L^{2}-E^{2}\right)=0.$ (7)
Assuming that $\dot{\varphi}>0$, we divide Eq. (7) by $\dot{\varphi}^{2}$ to yield the radial orbital differential equation:
$\bigg{(}\frac{\mathrm{d}\rho}{\mathrm{d}\varphi}\bigg{)}^{2}+V(\rho)=0,$ (8)
where the effective potential $V(\rho)$ is defined as
$V(\rho)=e^{-2\>\!\gamma}W^{2}\left(1-\frac{e^{-4\>\!\psi}W^{2}}{b^{2}}\right),$ (9)
with the impact parameter defined as $b\equiv L/E$.
Next, we consider circular orbits, for which $\dot{\rho}=0$ and $\ddot{\rho}=0$. Consequently, the effective potential $V$ and its derivative $V^{\prime}$, vanish at $\rho=\rho_{\mathrm{m}}$, where $\rho_{\mathrm{m}}$ denotes the coordinate radius of the circular orbit. Setting $V(\rho_{\mathrm{m}})=0$ immediately yields the critical impact parameter $b=b_{\mathrm{c}}$, with
$b_{\mathrm{c}}\equiv e^{-2\>\!\psi_{\mathrm{m}}}W_{\mathrm{m}},$ (10)
where, hereafter, the subscript $\mathrm{m}$ indicates evaluation at $\rho=\rho_{\mathrm{m}}$ and $\zeta=0$ [e.g., $\psi_{\mathrm{m}}\equiv\psi(\rho_{\mathrm{m}},0)$ and $W_{\mathrm{m}}\equiv W(\rho_{\mathrm{m}},0)$]. Furthermore, imposing $V^{\prime}(\rho_{\mathrm{m}})=0$ leads to
$\psi^{\prime}_{\mathrm{m}}=\frac{W^{\prime}_{\mathrm{m}}}{2W_{\mathrm{m}}}.$ (11)
Evaluating $V^{\prime\prime}$ at $\rho=\rho_{\mathrm{m}}$ for $b=b_{\mathrm{c}}$, we obtain
$V^{\prime\prime}_{\mathrm{m}}=2e^{-2\gamma_{\mathrm{m}}}\left[\>\!2W_{\mathrm{m}}^{2}\psi^{\prime\prime}_{\mathrm{m}}+(W^{\prime}_{\mathrm{m}})^{2}-W_{\mathrm{m}}W^{\prime\prime}_{\mathrm{m}}\>\!\right],$ (12)
which provides a criterion for the stability of the circular orbits: the orbit is unstable if $V^{\prime\prime}_{\mathrm{m}}<0$, and stable if $V^{\prime\prime}_{\mathrm{m}}>0$. In what follows, we focus on the unstable photon circular orbits (i.e., $V^{\prime\prime}_{\mathrm{m}}<0$), and hence,
$0\displaystyle 2W_{\mathrm{m}}^{2}\psi^{\prime\prime}_{\mathrm{m}}+(W^{\prime}_{\mathrm{m}})^{2}-W_{\mathrm{m}}W^{\prime\prime}_{\mathrm{m}}<$ 0 . (13)
III Deflection angle in the strong deflection limit
In this section, we derive the deflection angle in the SDL, following the approach in Ref. [4]. Let $\rho_{0}$ denote the radial coordinate of the closest approach, where $V(\rho_{0})=0$. Then, the impact parameter is given by
$b=e^{-2\>\!\psi_{0}}W_{0}.$ (14)
Throughout the manuscript, the subscript 00 denotes evaluation at $\rho=\rho_{0}$ and $\zeta=0$ [e.g., $\psi_{0}\equiv\psi(\rho_{0},0)$ and $W_{0}\equiv W(\rho_{0},0)$]. We define the integral
$I(\rho_{0})=2\int_{\rho_{0}}^{\rho_{\infty}}\frac{|\mathrm{d}\rho|}{\sqrt{-V}}$ (15)
which represents the total angular change experienced by a photon traveling from infinity to $\rho_{0}$ and then back to infinity. Finally, the deflection angle is defined by
$\alpha(\rho_{0})=I(\rho_{0})-\pi.$ (16)
To analyze the behavior of the deflection angle in the SDL, we introduce a new variable, as proposed in Ref. [35], defined by
$z=1-\frac{R_{0}}{R}.$ (17)
In terms of this variable, the integral $I(\rho_{0})$ becomes
$I(\rho_{0})=2\int_{0}^{1}\frac{\mathrm{d}z}{\sqrt{-\frac{(R^{\prime})^{2}}{R_{0}^{2}}(1-z)^{4}\>\!V}},$ (18)
where $R^{\prime}$ denotes the derivative of $R(\rho,0)$ with respect to $\rho$. By expanding $(R^{\prime})^{2}$ and $V$, as given in Eqs. (3) and (9), in powers of $z$, we obtain
$(R^{\prime})^{2}$ $=(R^{\prime}_{0})^{2}+2R_{0}R_{0}^{\prime\prime}z+O(z^{2}),$ (19)
$V$ $=\frac{R_{0}V^{\prime}_{0}}{R_{0}^{\prime}}\>\!z+\left[\>\!\left(\frac{R_{0}}{R_{0}^{\prime}}-\frac{R_{0}^{2}R_{0}^{\prime\prime}}{2\>\!(R_{0}^{\prime})^{3}}\right)V_{0}^{\prime}+\frac{R_{0}^{2}}{2\>\!(R_{0}^{\prime})^{2}}V_{0}^{\prime\prime}\>\!\right]z^{2}+O(z^{3}),$ (20)
Hence, truncating the expression under the square root in Eq. (18) to second order in $z$ yields $\sqrt{c_{1}z+c_{2}z^{2}}$. Accordingly, we define the corresponding integral as
$I_{\mathrm{D}}(\rho_{0})=2\int_{0}^{1}\frac{\mathrm{d}z}{\sqrt{c_{1}z+c_{2}z^{2}}},$ (21)
where the coefficients are given by
$c_{1}$ $=-\frac{R^{\prime}_{0}V^{\prime}_{0}}{R_{0}},$ (22)
$c_{2}$ $=3\left(\frac{R^{\prime}_{0}V^{\prime}_{0}}{R_{0}}-\frac{R^{\prime\prime}_{0}V^{\prime}_{0}}{2R^{\prime}_{0}}\right)-\frac{V^{\prime\prime}_{0}}{2}.$ (23)
The integral $I_{\mathrm{D}}$ isolates the leading-order divergent behavior of the total integral $I$ in the SDL, i.e., as $\rho_{0}\to\rho_{\mathrm{m}}$. Since $V^{\prime}_{0}\to 0$ in this limit, the integrand in Eq. (18) reduces to $1/(\sqrt{c_{2}}z)$, which leads to a logarithmic divergence. Thus, we identify $I_{\mathrm{D}}$ as the divergent part of $I$, and define the regular part as
$I_{\mathrm{R}}(\rho_{0})=I(\rho_{0})-I_{\mathrm{D}}(\rho_{0}).$ (24)
The regular part $I_{\mathrm{R}}$ typically may require numerical evaluation depending on the global structure of the spacetime. Evaluating the integral (21) yields
$I_{\mathrm{D}}(\rho_{0})=\frac{4}{\sqrt{c_{2}}}\log\frac{\sqrt{c_{1}+c_{2}}+\sqrt{c_{2}}}{\sqrt{c_{1}}}.$ (25)
To express the SDL (i.e., $\rho_{0}\to\rho_{\mathrm{m}}$) in a coordinate-independent manner, we adopt the impact parameter $b$ as a natural measure. This approach directly relates the deviation of $b$ from its critical value $b_{\mathrm{c}}$ to the small difference $(\rho_{0}-\rho_{\mathrm{m}})$. In particular, expanding $b$ around $b_{\mathrm{c}}$ yields
$b=b_{\mathrm{c}}\left[\>\!1-\frac{e^{2\>\!\gamma_{\mathrm{m}}}V^{\prime\prime}_{\mathrm{m}}}{4W_{\mathrm{m}}^{2}}(\rho_{0}-\rho_{\mathrm{m}})^{2}+O\left(\Big{(}\frac{\rho_{0}}{\rho_{\mathrm{m}}}-1\Big{)}^{3}\right)\>\!\right].$ (26)
Now, we assume that $R^{\prime}_{\mathrm{m}}\neq 0$ in what follows (see Appendix A for the case $R^{\prime}_{\mathrm{m}}=0$). Similarly, the coefficients $c_{1}$ and $c_{2}$ can be expanded in terms of $(\rho_{0}-\rho_{\mathrm{m}})$ as follows:
$c_{1}$ $=-\frac{R^{\prime}_{\mathrm{m}}V^{\prime\prime}_{\mathrm{m}}}{R_{\mathrm{m}}}(\rho_{0}-\rho_{\mathrm{m}})+O\left(\Big{(}\frac{\rho_{0}}{\rho_{\mathrm{m}}}-1\Big{)}^{2}\right),$ (27)
$c_{2}$ $=-\frac{V^{\prime\prime}_{\mathrm{m}}}{2}+O\left(\frac{\rho_{0}}{\rho_{\mathrm{m}}}-1\right).$ (28)
Alternatively, by inverting Eq. (26), we can express Eqs. (27) and (147) as
$c_{1}$ $=2\>\!e^{\psi_{\mathrm{m}}-\gamma_{\mathrm{m}}}R^{\prime}_{\mathrm{m}}\sqrt{-V^{\prime\prime}_{\mathrm{m}}}\left(\frac{b}{b_{\mathrm{c}}}-1\right)^{1/2}+O\left(\frac{b}{b_{\mathrm{c}}}-1\right),$ (29)
$c_{2}$ $=-\frac{V^{\prime\prime}_{\mathrm{m}}}{2}+O\left(\Big{(}\frac{b}{b_{\mathrm{c}}}-1\Big{)}^{1/2}\right).$ (30)
Using Eqs. (29) and (30), we can expand Eq. (25) in terms of $\left(b/b_{\mathrm{c}}-1\right)$ as
$I_{\mathrm{D}}(\rho_{0})=$ $-\sqrt{-\frac{2}{V^{\prime\prime}_{\mathrm{m}}}}\log\left(\frac{b}{b_{\mathrm{c}}}-1\right)$ (31)
$+\sqrt{-\frac{2}{V^{\prime\prime}_{\mathrm{m}}}}\log\left(-\frac{e^{2(\gamma_{\mathrm{m}}-\psi_{\mathrm{m}})}}{(R^{\prime}_{\mathrm{m}})^{2}}V^{\prime\prime}_{\mathrm{m}}\right)+O\left(\left(\frac{b}{b_{\mathrm{c}}}-1\right)^{1/2}\log\left(\frac{b}{b_{\mathrm{c}}}-1\right)\right).$ (32)
This result shows that the leading divergence is logarithmic. Consequently, the deflection angle in the SDL is given by
$\alpha(\rho_{0})=-\bar{a}\log\left(\frac{b}{b_{\mathrm{c}}}-1\right)+\bar{b}+O\left(\left(\frac{b}{b_{\mathrm{c}}}-1\right)^{1/2}\log\left(\frac{b}{b_{\mathrm{c}}}-1\right)\right),$ (33)
where the SDL coefficients $\bar{a}$ and $\bar{b}$ are defined as
$\bar{a}$ $=\sqrt{-\frac{2}{V^{\prime\prime}_{\mathrm{m}}}},$ (34)
$\bar{b}$ $=\bar{a}\log\left(\frac{2\>\!e^{2(\gamma_{\mathrm{m}}-\psi_{\mathrm{m}})}}{\bar{a}^{2}(R^{\prime}_{\mathrm{m}})^{2}}\right)+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi.$ (35)
Here, $\bar{a}$ quantifies the strength of the logarithmic divergence, while $\bar{b}$ represents the constant offset, i.e., the regular part of the deflection angle after subtracting the logarithmic divergence. These expressions are fundamental to determining the deflection angle in the SDL in a coordinate-invariant manner.
IV Coordinate-invariant form of the strong deflection coefficients via local curvature
In this section, we formulate the deflection angle in the SDL (33)–(35) using local and coordinate-invariant geometric quantities. To this end, we introduce the following tetrad:
$e_{(0)}^{a}$ $=e^{-\psi}\>\!(\partial/\partial t)^{a},$ (36)
$e_{(1)}^{a}$ $=e^{\psi-\gamma}\>\!(\partial/\partial\rho)^{a},$ (37)
$e_{(2)}^{a}$ $=e^{\psi-\gamma}\>\!(\partial/\partial\zeta)^{a},$ (38)
$e_{(3)}^{a}$ $=\frac{e^{\psi}}{W}\>\!(\partial/\partial\varphi)^{a}.$ (39)
Here, $e_{(0)}^{a}$ represents the four-velocity of static observers. The tetrad components of the Einstein tensor, $G_{(\mu)(\nu)}=G_{ab}e_{(\mu)}^{a}e_{(\nu)}^{b}$, have the following nonzero components:
$G_{(0)(0)}$ $=e^{2\>\!(\psi-\gamma)}\left(2\>\!(\psi_{\rho\rho}+\psi_{\zeta\zeta})-\psi_{\rho}^{2}-\psi_{\zeta}^{2}-\gamma_{\rho\rho}-\gamma_{\zeta\zeta}+\frac{2(\psi_{\rho}W_{\rho}+\psi_{\zeta}W_{\zeta})}{W}-\frac{W_{\rho\rho}+W_{\zeta\zeta}}{W}\right),$ (40)
$G_{(1)(1)}$ $=e^{2\>\!(\psi-\gamma)}\left(-\psi_{\rho}^{2}+\psi_{\zeta}^{2}+\frac{\gamma_{\rho}W_{\rho}-\gamma_{\zeta}W_{\zeta}}{W}+\frac{W_{\zeta\zeta}}{W}\right),$ (41)
$G_{(2)(2)}$ $=e^{2\>\!(\psi-\gamma)}\left(\psi_{\rho}^{2}-\psi_{\zeta}^{2}-\frac{\gamma_{\rho}W_{\rho}-\gamma_{\zeta}W_{\zeta}}{W}+\frac{W_{\rho\rho}}{W}\right),$ (42)
$G_{(1)(2)}$ $=e^{2\>\!(\psi-\gamma)}\left(-2\>\!\psi_{\rho}\psi_{\zeta}+\frac{\gamma_{\rho}W_{\zeta}+\gamma_{\zeta}W_{\rho}}{W}-\frac{W_{\rho\zeta}}{W}\right),$ (43)
$G_{(3)(3)}$ $=e^{2\>\!(\psi-\gamma)}\left(\psi_{\rho}^{2}+\psi_{\zeta}^{2}+\gamma_{\rho\rho}+\gamma_{\zeta\zeta}\right),$ (44)
with subscripts indicating partial differentiation (e.g., $\psi_{\rho}\equiv\partial\psi/\partial\rho$ and $\psi_{\rho\rho}\equiv\partial^{2}\psi/\partial\rho^{2}$).
Let $C_{abcd}$ denote the Weyl tensor, which encodes the free gravitational field in the spacetime. The electric part of the Weyl tensor with respect to the four-velocity $e_{(0)}^{a}$ (see, e.g., Ref. [45]) is defined by
$E_{ab}=C_{acbd}\>\!e_{(0)}^{c}e_{(0)}^{d}.$ (45)
The nonzero tetrad components of $E_{ab}$, defined by $E_{(\mu)(\nu)}=E_{ab}e_{(\mu)}^{a}e_{(\nu)}^{b}$, are given by
$E_{(1)(1)}=\frac{e^{2\>\!(\psi-\gamma)}}{6}\bigg{[}\>\!$ $4\>\!\psi_{\rho\rho}-2\>\!\psi_{\zeta\zeta}+8\>\!\psi_{\rho}^{2}-4\>\!\psi_{\zeta}^{2}-\gamma_{\rho\rho}-\gamma_{\zeta\zeta}-6\left(\psi_{\rho}\gamma_{\rho}-\psi_{\zeta}\gamma_{\zeta}\right)$ (46)
$-\frac{W_{\rho}\left(2\>\!\psi_{\rho}-3\>\!\gamma_{\rho}\right)}{W}-\frac{W_{\zeta}\left(2\>\!\psi_{\zeta}+3\>\!\gamma_{\zeta}\right)}{W}-\frac{W_{\rho\rho}-2W_{\zeta\zeta}}{W}\>\!\bigg{]},$ (47)
$E_{(2)(2)}=\frac{e^{2\>\!(\psi-\gamma)}}{6}\bigg{[}\>\!$ $4\>\!\psi_{\zeta\zeta}-2\>\!\psi_{\rho\rho}+8\>\!\psi_{\zeta}^{2}-4\>\!\psi_{\rho}^{2}-\gamma_{\rho\rho}-\gamma_{\zeta\zeta}+6\left(\psi_{\rho}\gamma_{\rho}-\psi_{\zeta}\gamma_{\zeta}\right)$ (48)
$-\frac{W_{\rho}\left(2\>\!\psi_{\rho}+3\>\!\gamma_{\rho}\right)}{W}-\frac{W_{\zeta}\left(2\>\!\psi_{\zeta}-3\>\!\gamma_{\zeta}\right)}{W}-\frac{W_{\zeta\zeta}-2\>\!W_{\rho\rho}}{W}\>\!\bigg{]},$ (49)
$E_{(3)(3)}=\frac{e^{2\>\!(\psi-\gamma)}}{6}\bigg{[}\>\!$ $-2\left(\psi_{\rho\rho}+\psi_{\zeta\zeta}\right)-4\left(\psi_{\rho}^{2}+\psi_{\zeta}^{2}\right)+2\left(\gamma_{\rho\rho}+\gamma_{\zeta\zeta}\right)$ (50)
$+\frac{4\left(\psi_{\rho}W_{\rho}+\psi_{\zeta}W_{\zeta}\right)}{W}-\frac{W_{\rho\rho}+W_{\zeta\zeta}}{W}\>\!\bigg{]},$ (51)
$E_{(1)(2)}=E_{(2)(1)}=e^{2\>\!(\psi-\gamma)}\bigg{[}\>\!$ $\psi_{\rho\>\!\zeta}+2\psi_{\rho}\psi_{\zeta}-\psi_{\rho}\gamma_{\zeta}-\psi_{\zeta}\gamma_{\rho}+\frac{W_{\rho}\gamma_{\zeta}+W_{\zeta}\gamma_{\rho}-W_{\rho\>\!\zeta}}{2W}\>\!\bigg{]}.$ (52)
Since the spacetime is static, the magnetic part of $C_{abcd}$ identically vanishes. Moreover, because the Weyl tensor is trace-free, we have
$E_{(1)(1)}+E_{(2)(2)}+E_{(3)(3)}=0.$ (53)
Due to the $\mathbb{Z}_{2}$ symmetry about the $\zeta=0$ plane [see Eq. (2)], we obtain
$\psi_{\zeta}(\rho,0)=0,\quad\gamma_{\zeta}(\rho,0)=0,\quad W_{\zeta}(\rho,0)=0.$ (54)
Furthermore, any term involving an odd number of $\zeta$-derivatives (e.g., $\psi_{\rho\zeta}$) vanishes on the plane. Consequently, the off-diagonal components of both the electric part of the Weyl tensor and the Einstein tensor vanish on the plane (i.e., $E_{(1)(2)}=E_{(2)(1)}=0$ and $G_{(1)(2)}=0$ at $\zeta=0$).
Using these curvature quantities, the second derivative of the effective potential evaluated at the unstable photon circular orbit [see Eq. (12)] is given by
$V^{\prime\prime}_{\mathrm{m}}=-2R^{2}_{\mathrm{m}}\left[\>\!E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}-\frac{G_{(0)(0)}^{\mathrm{m}}+G_{(3)(3)}^{\mathrm{m}}}{2}\>\!\right],$ (55)
where the subscript $\mathrm{m}$ of $E_{(\mu)(\nu)}$ and $G_{(\mu)(\nu)}$ indicates evaluation at $\rho=\rho_{\mathrm{m}}$ and $\zeta=0$. For the photon circular orbit to be unstable, we require that $V^{\prime\prime}_{\mathrm{m}}<0$; hence,
$E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}>\frac{G_{(0)(0)}^{\mathrm{m}}+G_{(3)(3)}^{\mathrm{m}}}{2}.$ (56)
This inequality establishes a local, coordinate-invariant relation between the Weyl and Einstein tensors (through their tetrad components) at the unstable photon circular orbit (see also Ref. [46]). In the following section, we reinterpret this inequality in terms of the relationship between the matter field quantities and the tidal forces.
Finally, the SDL coefficients given in Eqs. (34) and (35) can be expressed in terms of curvature quantities as
$\bar{a}$ $=\frac{1}{R_{\mathrm{m}}\sqrt{E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}-\frac{1}{2}\>\!\big{(}G_{(0)(0)}^{\mathrm{m}}+G_{(3)(3)}^{\mathrm{m}}\big{)}}},$ (57)
$\bar{b}$ $=\bar{a}\log\left[\>\!\frac{12}{\bar{a}^{2}R_{\mathrm{m}}^{2}\left[\>\!6E_{(3)(3)}^{\mathrm{m}}+G_{(0)(0)}^{\mathrm{m}}-G_{(3)(3)}^{\mathrm{m}}+2\>\!G_{(1)(1)}^{\mathrm{m}}+2\>\!G_{(2)(2)}^{\mathrm{m}}\>\!\right]}\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi,$ (58)
where $R^{\prime}_{\mathrm{m}}\neq 0$. These expressions show that the contributions from $I_{\mathrm{D}}$ to the SDL coefficients depend solely on $R_{\mathrm{m}}$ and the local, coordinate-invariant curvature quantities, $E_{(i)(i)}^{\mathrm{m}}$ and $G_{(\mu)(\mu)}^{\mathrm{m}}$. A comparison with the static, spherically symmetric analysis [35] reveals a key difference: whereas the electric part of the Weyl tensor is absent in that case, it is essential in the present analysis. This underscores the influence of the tidal effects in more general spacetimes.
We note that, although the contribution to $\bar{b}$ arising from $I_{\mathrm{D}}$ is fully determined by local curvature quantities, the remaining part $I_{\mathrm{R}}$ may involve global information about the spacetime. In particular, the decomposition into divergent and regular parts is not unique, and depends on the choice of the variable $z$.
V Relation between strong deflection limit coefficients and matter field quantities
In this section, we relate the SDL coefficients to the matter field quantities. We define a tensor $T_{(\mu)(\nu)}$ via the relation
$G_{(\mu)(\nu)}=8\pi\>\!T_{(\mu)(\nu)},$ (59)
which, in the context of general relativity, represents the Einstein equations with $T_{(\mu)(\nu)}$ interpreted as the energy-momentum tensor. In alternative theories of gravity, however, $T_{(\mu)(\nu)}$ may be understood as a more general tensor that not only encodes the energy-momentum distribution of matter fields but also incorporates additional curvature contributions. Without assuming a fixed physical interpretation for $T_{(\mu)(\nu)}$, we establish a framework in which the SDL coefficients remain independent of any specific gravitational theory. For simplicity, we will henceforth refer to $T_{(\mu)(\nu)}$ as the matter field quantities.
Let $T_{(\mu)(\nu)}^{\mathrm{m}}$ denote the value of $T_{(\mu)(\nu)}$ evaluated at $\rho=\rho_{\mathrm{m}}$ on the $\zeta=0$ plane (i.e., at the unstable photon circular orbit). Then, $V^{\prime\prime}_{\mathrm{m}}$ given in Eq. (55) can be written as
$V^{\prime\prime}_{\mathrm{m}}=-2R^{2}_{\mathrm{m}}\left[\>\!E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}-4\pi\left(T_{(0)(0)}^{\mathrm{m}}+T_{(3)(3)}^{\mathrm{m}}\right)\>\!\right].$ (60)
For $V$ to exhibit a local maximum at $\rho=\rho_{\mathrm{m}}$ on the $\zeta=0$ plane, we require that $V^{\prime\prime}_{\mathrm{m}}<0$; equivalently,
$E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}>4\pi\left(T_{(0)(0)}^{\mathrm{m}}+T_{(3)(3)}^{\mathrm{m}}\right),$ (61)
as stated in Eq. (56). In other words, the existence of an unstable photon circular orbit requires that the difference between the tidal components, $E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}$, is sufficiently large compared to the net contribution from the local matter fields $T_{(0)(0)}^{\mathrm{m}}+T_{(3)(3)}^{\mathrm{m}}$. When the net matter contribution is negligible (e.g., tipically in vacuum of general relativity), the tidal effects alone determine the orbital instability as
$E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}>0.$ (62)
Conversely, if the matter contribution is too large relative to the tidal difference, the condition in Eq. (61) is violated, and the photon circular orbit becomes stable rather than unstable.
Finally, the SDL coefficients $\bar{a}$ and $\bar{b}$ can be expressed in terms of local matter field quantities as
$\bar{a}$ $=\frac{1}{R_{\mathrm{m}}\sqrt{E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}-4\pi\>\!\big{(}T_{(0)(0)}^{\mathrm{m}}+T_{(3)(3)}^{\mathrm{m}}\big{)}}},$ (63)
$\bar{b}$ $=\bar{a}\log\left[\>\!\frac{6}{\bar{a}^{2}R_{\mathrm{m}}^{2}\left[\>\!3E_{(3)(3)}^{\mathrm{m}}+4\pi\big{(}T_{(0)(0)}^{\mathrm{m}}-T_{(3)(3)}^{\mathrm{m}}+2\>\!T_{(1)(1)}^{\mathrm{m}}+2\>\!T_{(2)(2)}^{\mathrm{m}}\big{)}\>\!\right]}\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi,$ (64)
where $R^{\prime}_{\mathrm{m}}\neq 0$. These expressions reveal that the contributions from $I_{\mathrm{D}}$ to the SDL coefficients are determined locally by $R_{\mathrm{m}}$, $E_{(i)(i)}^{\mathrm{m}}$, and the matter field quantities $T_{(\mu)(\mu)}^{\mathrm{m}}$. In particular, the coefficient $\bar{a}$ is entirely governed by the balance between the tidal contribution, $R_{\mathrm{m}}^{2}(E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}})$, and the matter field contribution, $4\pi R_{\mathrm{m}}^{2}(T_{(0)(0)}^{\mathrm{m}}+T_{(3)(3)}^{\mathrm{m}})$. Similarly, the contribution to $\bar{b}$ associated with $I_{\mathrm{D}}$ depends on $\bar{a}$ and the local balance between the tidal effects, $3R_{\mathrm{m}}^{2}E_{(3)(3)}^{\mathrm{m}}$, and the matter effects, $4\pi R_{\mathrm{m}}^{2}(T_{(0)(0)}^{\mathrm{m}}-T_{(3)(3)}^{\mathrm{m}}+2\>\!T_{(1)(1)}^{\mathrm{m}}+2\>\!T_{(2)(2)}^{\mathrm{m}})$.
The coordinate-invariant expressions for the SDL coefficients derived above suggest a potential link between theoretical predictions and observable lensing features near compact objects. If $R_{\mathrm{m}}$ is known, the observationally inferred $\bar{a}$ directly reflects information about the local tidal structure and matter distribution. In combination with accurate modeling of the background geometry, even the subleading term $\bar{b}$ may provide additional insights into the curvature and matter profile near the unstable photon circular orbit.
We focus on the special case where the following relations are satisfied:
$T_{(1)(1)}^{\mathrm{m}}+T_{(2)(2)}^{\mathrm{m}}=T_{(3)(3)}^{\mathrm{m}}=-T_{(0)(0)}^{\mathrm{m}},$ (65)
which are trivially satisfied in the vacuum of general relativity. Under these conditions, the expressions (63) and (64) reduce to
$\bar{a}$ $=\frac{1}{R_{\mathrm{m}}\sqrt{E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}}}.$ (66)
$\bar{b}$ $=\bar{a}\log\left(\frac{2}{\bar{a}^{2}R_{\mathrm{m}}^{2}E_{(3)(3)}^{\mathrm{m}}}\right)+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi.$ (67)
These results indicate that, when the matter field quantities satisfy the balance condition given in Eq. (65), the SDL coefficients are determined solely by the free gravitational field.
VI Strong deflection limit coefficients and Newman–Penrose scalars
To clarify the connection between the SDL coefficients and the intrinsic curvature of spacetime, we recast these coefficients in terms of NP scalars. This coordinate-invariant formulation directly links the observable features of gravitational lensing in the strong field limit to the fundamental curvature components of spacetime.
We introduce a null tetrad constructed from the orthonormal basis given in Eqs. (36)–(39). Specifically, we define
$l^{a}$ $=\frac{1}{\sqrt{2}}\Bigl{(}e_{(0)}^{a}+e_{(3)}^{a}\Bigr{)},$ (68)
$n^{a}$ $=\frac{1}{\sqrt{2}}\Bigl{(}e_{(0)}^{a}-e_{(3)}^{a}\Bigr{)},$ (69)
$m^{a}$ $=\frac{1}{\sqrt{2}}\Bigl{(}e_{(1)}^{a}+i\>\!e_{(2)}^{a}\Bigr{)},$ (70)
$\bar{m}^{a}$ $=\frac{1}{\sqrt{2}}\Bigl{(}e_{(1)}^{a}-i\>\!e_{(2)}^{a}\Bigr{)},$ (71)
where $i$ denotes the imaginary unit. These null vectors satisfy the standard normalization conditions $l^{a}n_{a}=-1$ and $m^{a}\bar{m}_{a}=1$, with all other scalar products vanishing.
We now project curvature tensors onto this null tetrad to obtain the NP scalars. Following Ref. [45], we define the Weyl scalars as
$\Psi_{0}$ $=C_{abcd}\>\!l^{a}\>\!m^{b}\>\!l^{c}\>\!m^{d},$ (72)
$\Psi_{1}$ $=C_{abcd}\>\!l^{a}\>\!n^{b}\>\!l^{c}\>\!m^{d},$ (73)
$\Psi_{2}$ $=C_{abcd}\>\!l^{a}\>\!m^{b}\>\!\bar{m}^{c}\>\!n^{d},$ (74)
$\Psi_{3}$ $=C_{abcd}\>\!l^{a}\>\!n^{b}\>\!\bar{m}^{c}\>\!n^{d},$ (75)
$\Psi_{4}$ $=C_{abcd}\>\!\bar{m}^{a}\>\!n^{b}\>\!\bar{m}^{c}\>\!n^{d}.$ (76)
These scalars encode the free gravitational field in a coordinate-invariant manner. For a static spacetime, we find that $\Psi_{4}+\Psi_{0}$ is real, $\Psi_{4}-\Psi_{0}$ is purely imaginary, and $\Psi_{1}=\Psi_{3}=0$. Moreover, $\mathbb{Z}_{2}$ symmetry about the $\zeta=0$ plane enforces $\Psi_{0}=\Psi_{4}$, which implies that $\Psi_{0}$ is real on that plane. Under these conditions, the electric part of the Weyl tensor in the chosen tetrad basis are given by
$E_{(i)(j)}=\left[\>\!\begin{array}[]{ccc}\Psi_{4}-\Psi_{2}&0&0\\ 0&-\Psi_{4}-\Psi_{2}&0\\ 0&0&2\Psi_{2}\end{array}\>\!\right].$ (80)
We also introduce the NP-Ricci scalars as
$\Phi_{00}$ $=\frac{1}{2}R_{ab}\>\!l^{a}l^{b},$ (81)
$\Phi_{11}$ $=\frac{1}{4}R_{ab}(l^{a}n^{b}+m^{a}\bar{m}^{b}),$ (82)
$\Phi_{22}$ $=\frac{1}{2}R_{ab}n^{a}n^{b},$ (83)
$\Phi_{01}$ $=\frac{1}{2}R_{ab}\>\!l^{a}m^{b},$ (84)
$\Phi_{02}$ $=\frac{1}{2}R_{ab}\>\!m^{a}m^{b},$ (85)
$\Phi_{12}$ $=\frac{1}{2}R_{ab}\>\!m^{a}n^{b}.$ (86)
These scalars encode the local matter contributions via the Einstein equations. For a static spacetime, we find that $\Phi_{01}=0$, $\Phi_{12}=0$, and $\Phi_{00}=\Phi_{22}$. Moreover, $\mathbb{Z}_{2}$ symmetry about the $\zeta=0$ plane enforces that $\Phi_{02}$ is real on that plane. Under these conditions, the nonzero NP-Ricci scalars on the $\zeta=0$ plane are given by
$\Phi_{00}$ $=\Phi_{22}=\frac{G_{(0)(0)}+G_{(3)(3)}}{4},$ (87)
$\Phi_{11}$ $=\frac{G_{(0)(0)}-G_{(3)(3)}+G_{(1)(1)}+G_{(2)(2)}}{8},$ (88)
$\Phi_{02}$ $=\frac{G_{(1)(1)}-G_{(2)(2)}}{4}.$ (89)
We denote by $\Psi_{0}^{\mathrm{m}}$, $\Psi_{2}^{\mathrm{m}}$, and $\Psi_{4}^{\mathrm{m}}$ the Weyl scalars evaluated at $\rho=\rho_{\mathrm{m}}$ on the $\zeta=0$ plane (i.e., at the unstable photon circular orbit). Similarly, the NP-Ricci scalars at the same location are denoted by $\Phi_{00}^{\mathrm{m}}$, $\Phi_{11}^{\mathrm{m}}$, and $\Phi_{22}^{\mathrm{m}}$. Recalling Eqs. (55), we find that $V^{\prime\prime}_{\mathrm{m}}$ is given by
$V^{\prime\prime}_{\mathrm{m}}=4R_{\mathrm{m}}^{2}\left(\Psi_{0}^{\mathrm{m}}+\Phi_{00}^{\mathrm{m}}\right).$ (90)
Since the orbit is unstable (i.e., $V^{\prime\prime}_{\mathrm{m}}<0$ as required in Eq. (56)), it follows that
$\Psi_{0}^{\mathrm{m}}+\Phi_{00}^{\mathrm{m}}<0,$ (91)
which encapsulates the necessary balance between the free gravitational field—encoded in $\Psi_{0}^{\mathrm{m}}$ (or $\Psi_{4}^{\mathrm{m}}$)—and the local matter contributions—captured by $\Phi_{00}^{\mathrm{m}}$. Thus, the NP formalism greatly simplifies the criterion for the instability of photon circular orbits.
$\Psi_{0}^{\mathrm{m}}<0,$ (92)
Finally, we obtain the alternative forms of the SDL coefficients in terms of the NP scalars as
$\bar{a}$ $=\frac{1}{R_{\mathrm{m}}\sqrt{-2\big{(}\Psi_{0}^{\mathrm{m}}+\Phi_{00}^{\mathrm{m}}\big{)}}},$ (93)
$\bar{b}$ $=\bar{a}\log\left[\>\!\frac{1}{\bar{a}^{2}R_{\mathrm{m}}^{2}\big{(}\Psi_{2}^{\mathrm{m}}+\Phi_{11}^{\mathrm{m}}-\mathcal{R}^{\mathrm{m}}/24\big{)}}\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi.$ (94)
where $\mathcal{R}^{\mathrm{m}}$ denotes the Ricci scalar at $\rho=\rho_{\mathrm{m}}$ and $\zeta=0$. The coefficient $\bar{a}$ scales only with $R_{\mathrm{m}}^{2}(\Psi_{0}^{\mathrm{m}}+\Phi_{00}^{\mathrm{m}})$, revealing that the tidal effects are captured by $\Psi_{0}^{\mathrm{m}}$, while the matter contributions are reflected in $\Phi_{00}^{\mathrm{m}}$. Similarly, the contribution to $\bar{b}$ from $I_{\mathrm{D}}$ depends on $\bar{a}$ and on $R_{\mathrm{m}}^{2}(\Psi_{2}^{\mathrm{m}}+\Phi_{11}^{\mathrm{m}}-\mathcal{R}^{\mathrm{m}}/24)$. These expressions relate the observable deflection angle in the SDL to specific combinations of NP scalars in the strong gravitational field regime. In particular, if the circumferential radius $R_{\mathrm{m}}$ is known, the combination $\Psi_{0}^{\mathrm{m}}+\Phi_{00}^{\mathrm{m}}$ can be extracted from observational data. Furthermore, by adopting a specific model, we can estimate $\Psi_{2}^{\mathrm{m}}+\Phi_{11}^{\mathrm{m}}-\mathcal{R}^{\mathrm{m}}/24$ based on observations. This NP-based formulation thus provides a coordinate-invariant framework for connecting lensing observations with local spacetime curvature, and may serve as a practical tool for probing strong gravity regions near compact objects.
VII Evaluation of the formalism in specific geometries
We illustrate the consistency and physical relevance of the general formalism by evaluating the SDL coefficients in several explicit static, axisymmetric spacetimes, showing how local curvature and matter fields determine light propagation near unstable photon orbits.
VII.1 Zipoy–Voorhees spacetimes
We consider the Zipoy–Voorhees spacetimes, which satisfy the vacuum Einstein equations. The metric functions in Eq. (1) are given by
$e^{2\psi}$ $=\left(\frac{R_{+}+R_{-}-2\ell}{R_{+}+R_{-}+2\ell}\right)^{\delta},\quad e^{2\gamma}=\left(\frac{(R_{+}+R_{-})^{2}-4\ell^{2}}{4R_{+}R_{-}}\right)^{\delta^{2}},\quad W=\rho,$ (95)
with
$R_{\pm}=\sqrt{\rho^{2}+(\zeta\pm\ell)^{2}}.$ (96)
Here, $\ell$ and $m$ are positive constants, and the deformation parameter $\delta$ is defined by
$\delta=\frac{m}{\ell}.$ (97)
When $\delta=1$, this metric recovers to the Schwarzschild metric.
Solving Eq. (11) yields the coordinate radius of the photon circular orbit as
$\rho_{\mathrm{m}}=m\sqrt{4-\frac{1}{\delta^{2}}},$ (98)
which implies that $\rho_{\mathrm{m}}$ exists only for $\delta>1/2$. The corresponding circumferential radius of the photon circular orbit, $R_{\mathrm{m}}$, and the critical impact parameter, given in Eq. (10), are
$R_{\mathrm{m}}$ $=\left(2+\frac{1}{\delta}\right)\left(\frac{2\delta-1}{2\delta+1}\right)^{\frac{1-\delta}{2}}m,$ (99)
$b_{\mathrm{c}}$ $=\left(2+\frac{1}{\delta}\right)\left(\frac{2\delta-1}{2\delta+1}\right)^{\frac{1-2\delta}{2}}m.$ (100)
Note that these results satisfy the following relation:
$b_{\mathrm{c}}=\left(\frac{2\delta+1}{2\delta-1}\right)^{\frac{\delta}{2}}R_{\mathrm{m}}.$ (101)
When $\delta=1$ (i.e., the Schwarzschild case), $R_{\mathrm{m}}=3m$ and $b_{\mathrm{c}}=3\sqrt{3}m$. The second derivative of the effective potential at the photon circular orbit is given by
$V^{\prime\prime}_{\mathrm{m}}=-2R_{\mathrm{m}}^{2}(E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}})=-2\left(\frac{4\delta^{2}-1}{4\delta^{2}}\right)^{1-\delta^{2}}<0,$ (102)
which indicates that the photon circular orbit is unstable.
The tetrad components of the Einstein tensor identically vanish. The nontrivial components of $E_{(\mu)(\nu)}^{\mathrm{m}}$ are
$R_{\mathrm{m}}^{2}E_{(1)(1)}^{\mathrm{m}}$ $=-\frac{5\delta^{2}-1}{8\delta^{2}}\left(\frac{4\delta^{2}}{4\delta^{2}-1}\right)^{\!\delta^{2}},$ (103)
$R_{\mathrm{m}}^{2}E_{(2)(2)}^{\mathrm{m}}$ $=\frac{3\delta^{2}-1}{8\delta^{2}}\left(\frac{4\delta^{2}}{4\delta^{2}-1}\right)^{\!\delta^{2}},$ (104)
$R_{\mathrm{m}}^{2}E_{(3)(3)}^{\mathrm{m}}$ $=\frac{1}{4}\left(\frac{4\delta^{2}}{4\delta^{2}-1}\right)^{\delta^{2}}.$ (105)
Correspondingly, the nontrivial NP scalars are given by
$R_{\mathrm{m}}^{2}\Psi_{0}^{\mathrm{m}}$ $=R_{\mathrm{m}}^{2}\Psi_{4}^{\mathrm{m}}=-\frac{1}{2}\left(\frac{4\delta^{2}-1}{4\delta^{2}}\right)^{1-\delta^{2}},$ (106)
$R_{\mathrm{m}}^{2}\Psi_{2}^{\mathrm{m}}$ $=\frac{1}{8}\left(\frac{4\delta^{2}}{4\delta^{2}-1}\right)^{\delta^{2}},$ (107)
Finally, we obtain the explicit forms of Eqs. (66) and (67) as
$\bar{a}$ $=\left(\frac{4\delta^{2}}{4\delta^{2}-1}\right)^{\frac{1-\delta^{2}}{2}},$ (108)
$\bar{b}$ $=\bar{a}\log\left(2\>\!\frac{4\delta^{2}-1}{\delta^{2}}\right)+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi.$ (109)
The result of $\bar{a}$ coincides with that of Ref. [33], providing a concrete example that supports our main results. As mentioned in Sec. IV, the contribution from $I_{\mathrm{D}}$ to $\bar{b}$ depends on the choice of $z$.
In the static, spherically symmetric limit, we obtain $\bar{a}=1$ and $\bar{b}=\log 6+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi$ (see, e.g., Refs. [4, 5, 35]). This limiting case implies
$E_{(2)(2)}^{\mathrm{m}}=E_{(3)(3)}^{\mathrm{m}}=-\frac{E_{(1)(1)}^{\mathrm{m}}}{2}=\frac{1}{3R_{\mathrm{m}}^{2}}.$ (110)
Here, the equality $E_{(2)(2)}=E_{(3)(3)}$ follows directly from spherical symmetry, and the trace-free property of the Weyl tensor then yields $E_{(1)(1)}=-2E_{(2)(2)}$. This degeneracy in $E_{(i)(j)}$, together with the absence of matter fields, results in the universal value $\bar{a}=1$. Under this degeneracy, the NP scalars reduce to
$R_{\mathrm{m}}^{2}\Psi_{4}^{\mathrm{m}}=R_{\mathrm{m}}^{2}\Psi_{0}^{\mathrm{m}}=-\frac{1}{2},\quad R_{\mathrm{m}}^{2}\Psi_{2}^{\mathrm{m}}=\frac{1}{6}.$ (111)
VII.2 Reissner–Nordström spacetimes
We consider the Reissner–Nordström spacetimes characterized by mass $M$ and electric charge $Q$. In the metric form presented in Eq. (1), the metric functions are given by (see, e.g., Ref. [44])
$e^{2\>\!\psi}=\frac{(R_{+}+R_{-})^{2}-4d^{2}}{(R_{+}+R_{-}+2M)^{2}},\quad e^{2\>\!\gamma}=\frac{(R_{+}+R_{-})^{2}-4d^{2}}{4R_{+}R_{-}},\quad W=\rho,$ (112)
where $d=\sqrt{M^{2}-Q^{2}}$ and $R_{\pm}=\sqrt{\rho^{2}+(\zeta\pm d)^{2}}$. When $Q=0$, this metric reduces to the Schwarzschild metric.
Solving Eq. (11) yields the coordinate radius of the photon circular orbit:
$\rho_{\mathrm{m}}=\sqrt{(R_{\mathrm{m}}-M)^{2}-d^{2}},$ (113)
where $R_{\mathrm{m}}$ is given by
$R_{\mathrm{m}}=\frac{3M+\sqrt{9M^{2}-8Q^{2}}}{2},$ (114)
indicating that photon circular orbits exist only when $80\leq Q^{2}\leq 9M^{2}/$ 8. The critical impact parameter (10) is
$b_{\mathrm{c}}=R_{\mathrm{m}}\sqrt{\frac{3}{1-Q^{2}/R_{\mathrm{m}}^{2}}}.$ (115)
The second derivative of the effective potential at the photon circular orbit is given by
$V^{\prime\prime}_{\mathrm{m}}=-2\left(2-\frac{3M}{R_{\mathrm{m}}}\right).$ (116)
The condition $V^{\prime\prime}_{\mathrm{m}}<0$ further restricts the electric charge to the range $80\leq Q^{2}<9M^{2}/$ 8.
The nontrivial components of $T_{(\mu)(\nu)}^{\mathrm{m}}$ are given by
$T_{(0)(0)}^{\mathrm{m}}=T_{(2)(2)}^{\mathrm{m}}=T_{(3)(3)}^{\mathrm{m}}=-T_{(1)(1)}^{\mathrm{m}}=\frac{3M-R_{\mathrm{m}}}{16\pi R_{\mathrm{m}}^{3}}.$ (117)
From these expressions, the corresponding components of $G_{(\mu)(\mu)}^{\mathrm{m}}$ can be derived. The nontrivial components of $E_{(\mu)(\nu)}^{\mathrm{m}}$ are given by
$E_{(2)(2)}^{\mathrm{m}}=E_{(3)(3)}^{\mathrm{m}}=-\frac{E_{(1)(1)}^{\mathrm{m}}}{2}=\frac{R_{\mathrm{m}}-M}{2R_{\mathrm{m}}^{3}},$ (118)
which reflects a degeneracy of the Weyl tensor due to spherical symmetry. Correspondingly, the nontrivial NP scalars are given by
$\Psi_{4}^{\mathrm{m}}=\Psi_{0}^{\mathrm{m}}=-3\Psi_{2}^{\mathrm{m}}=-\frac{3(R_{\mathrm{m}}-M)}{4R_{\mathrm{m}}^{3}},$ (119)
and
$\Phi_{00}^{\mathrm{m}}=\Phi_{22}^{\mathrm{m}}=\frac{Q^{2}}{2R_{\mathrm{m}}^{4}},\quad\Phi_{11}^{\mathrm{m}}=0,\quad\mathcal{R}^{\mathrm{m}}=0.$ (120)
Finally, the coefficients $\bar{a}$ and $\bar{b}$ given by Eqs. (63) and (64), or equivalently, Eqs. (93) and (94), are
$\bar{a}$ $=\sqrt{\frac{R_{\mathrm{m}}}{2R_{\mathrm{m}}-3M}},$ (121)
$\bar{b}$ $=\bar{a}\log\left(8-\frac{4M}{R_{\mathrm{m}}-M}\right)+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi.$ (122)
These results coincide with those obtained in Refs. [9, 10, 35], providing a consistency check for both $\bar{a}$ and $\bar{b}$. When $Q=M$, the coefficients reduce to $\bar{a}=\sqrt{2}$ and $\bar{b}=\sqrt{2}\log 4++I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi$.
VII.3 Majumdar–Papapetrou dihole spacetimes
We consider the Majumdar–Papapetrou dihole spacetimes, in which two extremal Reissner–Nordström black holes of equal mass $M$ are held in static equilibrium at a separation $a$. In the metric form presented in Eq. (1), the metric functions are given by
$e^{2\psi}=\left(1+\frac{M}{R_{+}}+\frac{M}{R_{-}}\right)^{-2},\quad\gamma=0,\quad W=\rho,$ (123)
where $R_{\pm}=\sqrt{\rho^{2}+(\zeta\pm a)^{2}}$. When $a=0$, the solution reduces to a single extremal Reissner–Nordström black hole with total mass $2M$ and charge $2M$. In what follows, we set $M=1$.
Solving Eq. (11) yields three distinct solutions for the coordinate radius of the photon circular orbit on the plane $\zeta=0$. Here, we select the branch corresponding to the unstable orbit, which is given by [31, 48, 47]
$\rho_{\mathrm{m}}=\sqrt{\frac{4}{9}\left(1+2\cos\left[\>\!\frac{1}{3}\arccos\left(1-\frac{27a^{2}}{4}\right)\>\!\right]\right)^{2}-a^{2}}.$ (124)
The instability of this orbit is confirmed by
$V^{\prime\prime}_{\mathrm{m}}=2+\frac{8\rho_{\mathrm{m}}^{2}\left[\>\!2\rho_{\mathrm{m}}^{2}+(a^{2}-2\rho_{\mathrm{m}}^{2})\big{(}2+\sqrt{\rho_{\mathrm{m}}^{2}+a^{2}}\big{)}\>\!\right]}{\left(\rho_{\mathrm{m}}^{2}+a^{2}\right)^{2}\big{(}2+\sqrt{\rho_{\mathrm{m}}^{2}+a^{2}}\big{)}^{2}}<0,$ (125)
for $90\leq a<a_{\infty}\equiv 2\sqrt{6}/$ 9 [49]. The corresponding circumferential radius and the critical impact parameter are given by
$R_{\mathrm{m}}$ $=\sqrt{\rho_{\mathrm{m}}}\left(1+\frac{2}{\sqrt{\rho_{\mathrm{m}}+a^{2}}}\right),$ (126)
$b_{\mathrm{c}}$ $=\sqrt{\rho_{\mathrm{m}}}\left(1+\frac{2}{\sqrt{\rho_{\mathrm{m}}+a^{2}}}\right)^{2},$ (127)
The nontrivial components of $T_{(\mu)(\nu)}^{\mathrm{m}}$ are
$T_{(0)(0)}^{\mathrm{m}}=T_{(2)(2)}^{\mathrm{m}}=T_{(3)(3)}^{\mathrm{m}}=-T_{(1)(1)}^{\mathrm{m}}=\frac{\rho_{\mathrm{m}}^{2}}{\pi\left(\rho_{\mathrm{m}}^{2}+a^{2}\right)^{2}\big{(}2+\sqrt{\rho_{\mathrm{m}}^{2}+a^{2}}\big{)}^{2}}.$ (128)
From these expressions, the components of $G_{(\mu)(\mu)}^{\mathrm{m}}$ can be derived. The nontrivial components of $E_{(\mu)(\nu)}^{\mathrm{m}}$ are
$E_{(1)(1)}^{\mathrm{m}}$ $=\frac{2\rho_{\mathrm{m}}^{2}\left[\>\!2a^{2}+(a^{2}-2\rho_{\mathrm{m}}^{2})\sqrt{\rho_{\mathrm{m}}^{2}+a^{2}}\>\!\right]}{R_{\mathrm{m}}^{2}\left(\rho_{\mathrm{m}}^{2}+a^{2}\right)^{2}\big{(}2+\sqrt{\rho_{\mathrm{m}}^{2}+a^{2}}\big{)}^{2}},$ (129)
$E_{(2)(2)}^{\mathrm{m}}$ $=\frac{2\rho_{\mathrm{m}}^{2}\left[\>\!-4a^{2}+(\rho_{\mathrm{m}}^{2}-2a^{2})\sqrt{\rho_{\mathrm{m}}^{2}+a^{2}}\>\!\right]}{R_{\mathrm{m}}^{2}\left(\rho_{\mathrm{m}}^{2}+a^{2}\right)^{2}\big{(}2+\sqrt{\rho_{\mathrm{m}}^{2}+a^{2}}\big{)}^{2}},$ (130)
$E_{(3)(3)}^{\mathrm{m}}$ $=\frac{2\rho_{\mathrm{m}}^{2}\left[\>\!2a^{2}+(\rho_{\mathrm{m}}^{2}+a^{2})^{3/2}\>\!\right]}{R_{\mathrm{m}}^{2}\left(\rho_{\mathrm{m}}^{2}+a^{2}\right)^{2}\big{(}2+\sqrt{\rho_{\mathrm{m}}^{2}+a^{2}}\big{)}^{2}}.$ (131)
Correspondingly, the nontrivial NP scalars are given by
$\Psi_{4}^{\mathrm{m}}$ $=\Psi_{0}^{\mathrm{m}}=\frac{3\rho_{\mathrm{m}}^{2}\left[\>\!a^{4}-\rho_{\mathrm{m}}^{4}+2a^{2}\sqrt{\rho_{\mathrm{m}}^{2}+a^{2}}\>\!\right]}{R_{\mathrm{m}}^{2}(\rho_{\mathrm{m}}^{2}+a^{2})^{5/2}\big{(}2+\sqrt{\rho_{\mathrm{m}}^{2}+a^{2}}\big{)}^{2}},$ (132)
$\Psi_{2}^{\mathrm{m}}$ $=\frac{\rho_{\mathrm{m}}^{2}\left[\>\!2a^{2}+(\rho_{\mathrm{m}}^{2}+a^{2})^{3/2}\>\!\right]}{R_{\mathrm{m}}^{2}(\rho_{\mathrm{m}}^{2}+a^{2})^{2}\big{(}2+\sqrt{\rho_{\mathrm{m}}^{2}+a^{2}}\big{)}^{2}},$ (133)
and
$\Phi_{00}^{\mathrm{m}}=\Phi_{22}^{\mathrm{m}}=-\Phi_{02}^{\mathrm{m}}=\frac{2\rho_{\mathrm{m}}^{2}}{R_{\mathrm{m}}^{2}\left(\rho_{\mathrm{m}}^{2}+a^{2}\right)^{2}\big{(}2+\sqrt{\rho_{\mathrm{m}}^{2}+a^{2}}\big{)}^{2}},\quad\Phi_{11}^{\mathrm{m}}=0,\quad\mathcal{R}^{\mathrm{m}}=0.$ (134)
Finally, the coefficients $\bar{a}$ and $\bar{b}$ given by Eqs. (63) and (64), or equivalently, Eqs. (93) and (94), are
$\bar{a}$ $=\left[\>\!-1-\frac{4\rho_{\mathrm{m}}^{2}\left[\>\!2\rho_{\mathrm{m}}^{2}+(a^{2}-2\rho_{\mathrm{m}}^{2})\big{(}2+\sqrt{\rho_{\mathrm{m}}^{2}+a^{2}}\big{)}\>\!\right]}{\left(\rho_{\mathrm{m}}^{2}+a^{2}\right)^{2}\big{(}2+\sqrt{\rho_{\mathrm{m}}^{2}+a^{2}}\big{)}^{2}}\>\!\right]^{-1/2},$ (135)
$\bar{b}$ $=\bar{a}\log\left[\>\!\frac{2\rho_{\mathrm{m}}^{2}\left(\rho_{\mathrm{m}}^{2}+a^{2}\right)\big{(}2+\sqrt{\rho_{\mathrm{m}}^{2}+a^{2}}\big{)}^{4}}{\bar{a}^{2}R_{\mathrm{m}}^{2}\big{(}2a^{2}+(\rho_{\mathrm{m}}^{2}+a^{2})^{3/2}\big{)}^{2}}\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi.$ (136)
We find that $\bar{a}$ coincides with the result in Ref. [31], providing a concrete example that supports our main results. As mentioned in Sec. IV, the contribution from $I_{\mathrm{D}}$ to $\bar{b}$ depends on the choice of $z$. When $a=0$, these coefficients reduce to those for $Q=M$ in Sec. VII.2.
VIII Strong deflection limit and quasinormal modes
A well-established connection exists between QNM frequencies and the instability of photon circular orbits in spherically symmetric spacetimes, particularly in the eikonal limit [36, 37]. The real part of the QNM frequency corresponds to the angular frequency of the photon circular orbit $\Omega_{\mathrm{c}}=1/b_{\mathrm{c}}$, while the imaginary part is determined by the Lyapunov exponent $\lambda_{\mathrm{L}}$ characterizing the instability of the orbit
$\omega_{\mathrm{QNM}}=\Omega_{\mathrm{c}}\>\!l-i\left(n+\frac{1}{2}\right)\lambda_{\mathrm{L}},$ (137)
where $n$ is the overtone number, $l$ is the angular momentum of the perturbation [37]. The Lyapunov exponent is given by (see also Refs.  [50, 51, 52] for recent topics)
$\lambda_{\mathrm{L}}=\frac{1}{b_{\mathrm{c}}}\sqrt{-\frac{V^{\prime\prime}_{\mathrm{m}}}{2}}.$ (138)
Note that this expression is rewritten using the present notation, for consistency with our SDL formalism, rather than the original expression in Ref. [37]. Several studies have further explored a potential correspondence between QNM parameters and the SDL coefficient that governs the logarithmic divergence rate [38, 39]
$\lambda_{\mathrm{L}}=\frac{1}{b_{\mathrm{c}}\bar{a}},$ (139)
which indicates that the Lyapunov exponent appearing in the QNM frequency expression is proportional to the inverse of the SDL coefficient $\bar{a}$. Notably, a similar approach has also been employed on the reflection-symmetric plane of axisymmetric spacetimes, where QNM frequencies are computed using equatorial photon circular orbits. These results suggest a deep geometric link between wave dynamics and lensing in strong gravity.
Recently, an explicit expression for $\lambda_{\mathrm{L}}$ in spherically symmetric spacetimes has been derived in terms of local geometric and matter field quantities [35]:
$\lambda_{\mathrm{L}}=\frac{\sqrt{1-8\pi R_{\mathrm{m}}^{2}(\rho_{\mathrm{m}}+\Pi_{\mathrm{m}})}}{b_{\mathrm{c}}}.$ (140)
Here, $\rho_{\mathrm{m}}$ and $\Pi_{\mathrm{m}}$ denote energy density and tangential pressure evaluated at the photon sphere with the areal radius $R_{\mathrm{m}}$. This result reveals the relation between the damping rate of QNMs and the local matter field quantities.
Our results in the present paper extend the formula (140) to static, axisymmetric spacetimes. Using Eqs. (63) or (93), we recast $\lambda_{\mathrm{L}}$ in the form of Eq. (139) as
$\lambda_{\mathrm{L}}=\frac{R_{\mathrm{m}}}{b_{\mathrm{c}}}\sqrt{E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}-4\pi\>\!\big{(}T_{(0)(0)}^{\mathrm{m}}+T_{(3)(3)}^{\mathrm{m}}\big{)}},$ (141)
or equivalently,
$\lambda_{\mathrm{L}}=\frac{R_{\mathrm{m}}\sqrt{-2\big{(}\Psi_{0}^{\mathrm{m}}+\Phi_{00}^{\mathrm{m}}\big{)}}}{b_{\mathrm{c}}}.$ (142)
These expressions are determined by the coordinate-independent, local curvature and matter field quantities evaluated at the unstable photon circular orbit. This indicates that the damping rate of QNMs is directly linked to the local geometric and matter field properties at the radius of the unstable photon circular orbit.
In the following, we test our geometric expression for the Lyapunov exponent in several static, axisymmetric spacetimes. For each case, we evaluate the NP scalars at the photon circular orbit and confirm agreement with the result obtained from the effective potential. Explicit computations are given in Sec. VII.
In the case of the Zipoy–Voorhees spacetimes in Sec. VII.1, where $\Phi_{00}^{\mathrm{m}}=0$, substituting Eqs. (99), (100), and (106) into the formula (142), we obtain
$\lambda_{\mathrm{L}}=\frac{1}{2m\delta}\left(\frac{4\delta^{2}-1}{4\delta^{2}}\right)\left(\frac{2\delta-1}{2\delta+1}\right)^{\delta}.$ (143)
Expanding around $\delta=1$, we obtain
$\lambda_{\mathrm{L}}=\frac{1}{3\sqrt{3}m}\left[\>\!1+\left(2\log\frac{2}{3}\right)(\delta-1)+O((\delta-1)^{2})\>\!\right],$ (144)
which agrees with the result obtained in Ref. [53]. The leading term $1/(3\sqrt{3}m)$ corresponds to the value for the Schwarzschild case.
In the case of the Reissner–Nordström spacetime in Sec. VII.2, substituting Eqs. (114), (115), and (119) into the formula (142), we obtain
$\lambda_{\mathrm{L}}=\frac{1}{R_{\mathrm{m}}}\sqrt{\left(\frac{2}{3}-\frac{M}{R_{\mathrm{m}}}\right)\left(1-\frac{Q^{2}}{R_{\mathrm{m}}^{2}}\right)},$ (145)
which agrees with the result obtained in Ref. [54]. When $Q=M$, the expression reduces to $\lambda_{\mathrm{L}}=1/(4\sqrt{2}M)$.
In the case of the Majumdar–Papapetrou dihole spacetimes in Sec. VII.3, substituting Eqs. (126), (127), (132), and (134) into the formula (142), we obtain
$\lambda_{\mathrm{L}}=\frac{\sqrt{d_{\mathrm{m}}^{4}(4+4d_{\mathrm{m}}-d_{\mathrm{m}}^{2})-4a^{2}d_{\mathrm{m}}(6+5d_{\mathrm{m}})+4a^{4}(4+3d_{\mathrm{m}})}}{(2+d_{\mathrm{m}})^{3}\sqrt{d_{\mathrm{m}}^{2}-a^{2}}},$ (146)
which agrees with the result obtained in Ref. [47]. When $a=0$, the expression reduces to $\lambda_{\mathrm{L}}=1/(8\sqrt{2})$.
These consistent results support the validity of the geometric formula (142) for the Lyapunov exponent in static, axisymmetric spacetimes.
IX Summary and discussion
We study the deflection angle of photons in the SDL in static, axisymmetric spacetimes with reflection symmetry across the equatorial plane. Following the standard method of isolating the logarithmic divergence, we introduce a new variable that allows for a coordinate-invariant formulation of the SDL. We show that the two SDL coefficients, which characterize the logarithmic divergence and its offset, are determined by the second derivative of the effective potential evaluated at the unstable photon circular orbit. This second derivative is further related to local curvature and matter field quantities, yielding a coordinate-invariant expression for these coefficients. This local curvature consists of two distinct contributions: the tidal effects associated with the free gravitational field, described by the electric part of the Weyl tensor, and the matter-induced curvature, encoded in the Einstein tensor, representing the influence of matter fields. We also recast the SDL coefficients in terms of NP scalars, offering a geometric perspective aligned with the null structure of the spacetime. This formulation suggests a promising framework for interpreting observational data in terms of fundamental curvature quantities, particularly through coordinate-invariant combinations of NP Weyl and Ricci scalars. We confirm the consistency of our formulation by applying it to several known spacetimes.
In static, spherically symmetric spacetimes, the SDL coefficients can be expressed solely in terms of the areal radius of the photon sphere and the local energy density and pressure [35]. In contrast, in static, axisymmetric spacetimes, the deflection angle in the SDL is governed not only by local matter fields but also by the free gravitational field encoded in tidal distortions, leading to a more intricate geometric structure. Nevertheless, the leading logarithmic divergence of the deflection angle is still determined entirely by local quantities with clear geometric meaning at the unstable photon circular orbit. This ensures that the formulation retains its coordinate-invariant character.
These results provide a geometrically transparent foundation for connecting theoretical predictions with observations. By expressing the SDL coefficients entirely in terms of coordinate-invariant local quantities, the present formalism establishes a clear link between observable lensing features and the underlying geometry and matter distribution near the unstable photon circular orbit. This, in turn, opens up the possibility of inferring local geometric or matter-field properties—such as tidal forces, energy density, and pressure—from precise measurements of strong gravitational lensing near compact objects.
Through the connection between the QNM and SDL, the present analysis expresses the QNM frequency in terms of local curvature and matter field quantities. This formulation opens up the possibility of probing the local matter and geometry near the unstable photon circular orbit through gravitational wave observations. Furthermore, our findings may also shed light on a universal upper bound on chaos in thermal quantum field theory [55, 56], which provides an inequality between the Lyapunov exponent and the surface gravity of the horizon, a relation whose generalization to the photon sphere has recently been discussed [50, 51].
The present formalism provides a practical framework for extracting physical information—such as energy density, pressure, and tidal structure—from the deflection angle near ultracompact objects. Applied to specific models, it may offer valuable insights into the physics of strong gravity and serve as a probe of alternative theories of gravity. A natural extension of this work is to generalize the formalism to stationary and axisymmetric spacetimes; this development is already underway and will be presented in a separate publication.
Acknowledgements.
The author gratefully acknowledges the useful discussions with Yohsuke Takamori. This work was supported in part by JSPS KAKENHI Grant Nos. JP22K03611, JP23KK0048, and JP24H00183 and by Gakushuin University.
Appendix A Analysis of the $R^{\prime}_{\mathrm{m}}=0$ case
We consider the case $R^{\prime}_{\mathrm{m}}=0$, which, for example, is encountered in wormhole geometries. Since this condition implies that $W^{\prime}_{\mathrm{m}}=W_{\mathrm{m}}\psi^{\prime}_{\mathrm{m}}$, combining it with Eq. (11) yields $W^{\prime}_{\mathrm{m}}=0$. In this case, the coefficients given in Eqs. (22) and (23) are expanded about $\rho_{0}=\rho_{\mathrm{m}}$ as
$c_{1}$ $=-\frac{R^{\prime\prime}_{\mathrm{m}}V^{\prime\prime}_{\mathrm{m}}}{R_{\mathrm{m}}}(\rho_{0}-\rho_{\mathrm{m}})^{2}+O\left(\Big{(}\frac{\rho_{0}}{\rho_{\mathrm{m}}}-1\Big{)}^{3}\right),$ (147)
$c_{2}$ $=-2V^{\prime\prime}_{\mathrm{m}}+\left(-\frac{3}{2}\frac{R^{\prime\prime\prime}_{\mathrm{m}}V^{\prime\prime}_{\mathrm{m}}}{R^{\prime\prime}_{\mathrm{m}}}-\frac{V^{\prime\prime\prime}_{\mathrm{m}}}{2}\right)(\rho_{0}-\rho_{\mathrm{m}})+O\left(\Big{(}\frac{\rho_{0}}{\rho_{\mathrm{m}}}-1\Big{)}^{2}\right).$ (148)
Inverting Eq. (26) allows us to rewrite Eqs. (147) and (148) in terms of $(b/b_{\mathrm{c}}-1)$ as
$c_{1}$ $=4R_{\mathrm{m}}R^{\prime\prime}_{\mathrm{m}}e^{2(\psi_{\mathrm{m}}-\gamma_{\mathrm{m}})}\left(\frac{b}{b_{\mathrm{c}}}-1\right)+O\left(\bigg{(}\frac{b}{b_{\mathrm{c}}}-1\bigg{)}^{3/2}\right),$ (149)
$c_{2}$ $=-2V^{\prime\prime}_{\mathrm{m}}-\frac{e^{\psi_{\mathrm{m}}-\gamma_{\mathrm{m}}}R_{\mathrm{m}}}{\sqrt{-V^{\prime\prime}_{\mathrm{m}}}}\left(\frac{3R^{\prime\prime\prime}_{\mathrm{m}}V^{\prime\prime}_{\mathrm{m}}}{R^{\prime\prime}_{\mathrm{m}}}+\frac{V^{\prime\prime\prime}_{\mathrm{m}}}{2}\right)\left(\frac{b}{b_{\mathrm{c}}}-1\right)^{1/2}+O\left(\frac{b}{b_{\mathrm{c}}}-1\right).$ (150)
Using Eqs. (25), (149), and (150), we obtain the deflection angle in the SDL (33), with the corresponding SDL coefficients given by
$\bar{a}$ $=\sqrt{-\frac{2}{V^{\prime\prime}_{\mathrm{m}}}},$ (151)
$\bar{b}$ $=\bar{a}\log\frac{4e^{2(\gamma_{\mathrm{m}}-\psi_{\mathrm{m}})}}{\bar{a}^{2}R_{\mathrm{m}}R^{\prime\prime}_{\mathrm{m}}}.$ (152)
Note that the coefficient $\bar{a}$ is identical to that obtained in the $R^{\prime}_{\mathrm{m}}\neq 0$ case; consequently, Eqs. (57), (63), and (93) remain unchanged. On the other hand, $\bar{b}$ is expressed in terms of the Einstein tensor as
$\bar{b}=\bar{a}\log\left[\>\!\frac{16}{\bar{a}^{2}R_{\mathrm{m}}^{2}\left[\>\!2(E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}})-G_{(0)(0)}^{\mathrm{m}}-G_{(3)(3)}^{\mathrm{m}}+2G_{(2)(2)}^{\mathrm{m}}\>\!\right]}\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi,$ (153)
in terms of matter field quantities as
$\bar{b}=\bar{a}\log\left[\>\!\frac{8}{\bar{a}^{2}R_{\mathrm{m}}^{2}\left[\>\!E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}-4\pi(T_{(0)(0)}^{\mathrm{m}}+T_{(3)(3)}^{\mathrm{m}}-2T_{(2)(2)}^{\mathrm{m}})\>\!\right]}\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi,$ (154)
and in terms of the NP scalars as
$\bar{b}=\bar{a}\log\left[\>\!\frac{8}{\bar{a}^{2}R_{\mathrm{m}}^{2}\left[\>\!-2(\Psi_{0}^{\mathrm{m}}+\Phi_{00}^{\mathrm{m}}-\Phi_{11}^{\mathrm{m}}+\Phi_{02}^{\mathrm{m}})-\mathcal{R}^{\mathrm{m}}/4\>\!\right]}\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi.$ (155)